OpenStudy (anonymous):

does anyone know where i can find the complete rules for the derivatives regarding e? I'm baffled by the first derivative of e^x...

OpenStudy (solomonzelman):

you want a proof ?

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle \frac{d}{dx}\left(e^x\right)=\lim_{h\rightarrow ~0}\frac{e^{x+h}-e^x}{h}}\)

OpenStudy (solomonzelman):

use the first principles of derivatives to prove it

OpenStudy (solomonzelman):

and, just to know \(\large\color{slate}{\displaystyle \frac{d}{dx}\left(e^x\right)=e^x}\)

OpenStudy (anonymous):

ohhhhhhhhhhhh

OpenStudy (solomonzelman):

which would not be exactly true with a different exponent

OpenStudy (anonymous):

thanks so much, wow this setup rules

OpenStudy (bibby):

I dunno if this is strictly "legal" but \(\dfrac{d}{dx} a^x = a^x ln(a)\)

OpenStudy (anonymous):

yep, bibby, now i remember.... i must have been looking at the wrong chapter in my notes

OpenStudy (solomonzelman):

or, if you are assuming that given that d/dx (1/x)=ln(x) then \(\large\color{slate}{\displaystyle y=e^x}\) \(\large\color{slate}{\displaystyle \ln(y)=\ln(e^x)}\) \(\large\color{slate}{\displaystyle \ln(y)=x\ln(e)}\) \(\large\color{slate}{\displaystyle \ln(y)=x}\) \(\large\color{slate}{\displaystyle y'~(1/y)=1}\) \(\large\color{slate}{\displaystyle y'=1\times y}\) \(\large\color{slate}{\displaystyle y'= y}\)\ and we know from the beginning y=e^x, so \(\large\color{slate}{\displaystyle y'=e^x}\)

OpenStudy (solomonzelman):

I mean given that d/dx lnx = 1/x the other way around... sorry

OpenStudy (bibby):

which means \(\dfrac{d}{dx}e^x=e^x*\ln(e)=e^x*1=e^x\)

OpenStudy (ybarrap):

When you take the Taylor series expansion of e^x you get $$ 1+x+x^2/2+x^3/6+x^4/24+x^5/120+x^6/720+\\ x^7/5040+x^8/40320+x^9/362880+x^{10}/3628800+O(x^{11}) $$ If you take the derivative of e^x, you can take the term by term derivative of the above, and again you get the Taylor Series expansion of e^x, which shows that the derivative of e^x is e^x

OpenStudy (solomonzelman):

yes, you would just be getting previosu term, and there are infinity of those

OpenStudy (solomonzelman):

very good approximation of e^x would it be.

OpenStudy (solomonzelman):

anyway...

OpenStudy (solomonzelman):

have fun, and keep in mind log. differentiation, if you ever forget or want to prove quotient or product rule for differentiating a function

OpenStudy (solomonzelman):

Oh, interesting excessive try showing it \(\large\color{slate}{\displaystyle \frac{d^n}{dx^n}\left(xe^x\right)=ne^x+xe^x}\)

OpenStudy (solomonzelman):

exercise (spelling check... )

OpenStudy (bibby):

can you use induction?

OpenStudy (solomonzelman):

well, I am talking about just taking derivative twice or 3 times and showing that it will be like this for any nth derivative

OpenStudy (bibby):

yeah, that's induction pretty much

OpenStudy (bibby):

as I remember it

OpenStudy (solomonzelman):

also derivative of cos(x) and sin(x) would be interesting.

OpenStudy (solomonzelman):

I mean if you use the first prinicples. all that matters is practice.... practice as much as you can and you will master it!

OpenStudy (anonymous):

woah, so is this Taylor expansion thing how you find an equation for which you have the derivative?

OpenStudy (ybarrap):

$$ {d\over dx}\left (1+x+x^2/2+x^3/6+x^4/24+x^5/120+x^6/720+\\ x^7/5040+x^8/40320+x^9/362880+x^{10}/3628800+O(x^{11})\right )\\ =1+x+x^2/2+x^3/6+x^4/24+x^5/120+x^6/720+\\ x^7/5040+x^8/40320+x^9/362880+x^{10}/3628800+O(x^{11}) $$